OpenStudy (anonymous):

can anybody solve partial deriavative z=e^x^2 + xy with respect to x and y

OpenStudy (yuki):

yes let's find f_x first

OpenStudy (anonymous):

When you take the partial derivative with respect to x, consider y as a constant.

OpenStudy (yuki):

\[f = e^{x^2} + xy \] \[f_x = 2x*e^{x^2} + y\]

OpenStudy (anonymous):

\[{\partial z \over \partial x}=2xe^{x^2}+y\]

OpenStudy (yuki):

\[f_y = 0 + x = x\]

OpenStudy (anonymous):

\[{\partial z \over \partial y}=x\]

OpenStudy (yuki):

Anwar you always surprise me. How did you get that "d" from ?!

OpenStudy (yuki):

Anwar you always surprise me. How did you get that "d" from ?!\[\partial\]

OpenStudy (anonymous):

Try to find out yourself :)

OpenStudy (anonymous):

cool !!!

OpenStudy (anonymous):

lol that was fast :)

OpenStudy (yuki):

that was a lot easier than I thought lol

OpenStudy (anonymous):

Haha yeah

OpenStudy (yuki):

anyway, zizUo, with partial derivatives you will treat the other variable as same as numbers, so for f_y, the term \[e^{x^2}\] has no y in it, so it's derivative is 0 since you treat is as if it's a number

OpenStudy (yuki):

so it is actually not that hard. implicit differentiation is much more harder :)

OpenStudy (anonymous):

it means u simply have to separate out the e^x ?

OpenStudy (yuki):

e^x^2 has no y in it, so the partial derivative of e^x^2 with respect to y, is as same as taking the derivative of a number like 10 or 34

OpenStudy (yuki):

on the other hand, the term xy has a y multiplied to x, so it is similar to saying " find the derivative of 2y" which is 2 in our case, x is the constant, so the partial derivative of xy is x

OpenStudy (yuki):

on the other hand, the term xy has a y multiplied to x, so it is similar to saying " find the derivative of 2y" which is 2 in our case, x is the constant, so the partial derivative of xy is x